Answer them as soon as possible please
D Question 1 3 pts Estimate the cost of developing a new clinic with 250000 ft2 space. The appropriate capacity exponent is 0.66. The cost of a clinic with 200000 ft2 was $20 million 5 years ago. The cost index was 200 five years ago, and 235 now. (a) What is the cost of a clinic with 200000 ft2 space now? O $21.3 million $23.5 million O $25.1 million O $29.0 million
D Question 2 3 pts (b) What is the cost of a clinic with 250000 ft2 space now? O $23.50 million $25.24 million O$27.23 million $29.65 million Quiz saved at 11:09pm Submit Quiz
Answer #1

1.

B

Working note:

Cost of clinic with 200000 square feet = 20*(235/200) = $23.5 Million

2.

C.

Working note:

Cost of clinic with 250000 square feet = 20*(235/200)*(250000/200000)^.66

Cost of clinic with 250000 square feet = $27.23 Million